OpenStudy (anonymous):

An astronaut at rest on Earth has a heartbeat rate of 74 beats/min. When the astronaut is traveling in a spaceship at 0.85c, what will this rate be as measured by the following.

OpenStudy (noelgreco):

the following:

OpenStudy (anonymous):

The astronaut will still measure his heart rate to be 74 beats per minute; however an observer on Earth would see the astronauts heart rate slow down. To those observers on Earth (or any observer stationary with respect to the astronaut's motion), the astronaut's heart rate would appear to be as follows: \[HR'=\frac{ HR }{ \gamma }=\left( \sqrt{1-\frac{ v ^{2} }{ c ^{2} }} \right)HR\]\[HR'=\left( \sqrt{1-0.86^{2}} \right)*74\]So HR'=37.8 beats per minute, the astronaut's apparent heart rate as measured by an observer.